If two sides of a square field were increased by five feet the area of the field would increase by 245ft2. Find the area of the original square.
Answer by checkley77(12844)

(X+5)^2=X^2+245
X^2+10X+25=X^2+245
X^2-X^2+10X=245-25
10X=220
X=220/10
X=22 THE ORIGINAL LENGTH OF THE SQUARE SIDES.
22+5=27 LENGTHS OF THE EXPANDED SQUARE SIDES.
PROOF:
27^2=22^2+245
729=484+245
729=729